X. 1 2 3 4 5


P(x) 1/6 a/6 5/24 a/6 a/12



Find a


Find P(4)


Find P(5)



Pa answer na po kaagad

1
Expert's answer
2022-04-11T10:14:25-0400

1/6+a/6+5/24+a/6+a/12=1


9/24+10a/24=1


10a/24=15/24


10a=15


a=1.5


P(4)=1.5/6=0.25


P(5)=1.5/12=0.125